OpenStudy (anonymous):

Suppose a real-world situation is modeled by a quadratic equation in the form of where t represents time. What must be true about the equation if the problem situation only has one solution?

OpenStudy (anonymous):

at to the power of 2 =bt+ c=0

OpenStudy (anonymous):

The comment i made is the form that the question is referring to.

OpenStudy (anonymous):

@Lily2913 can you help me or know anyone who can?

OpenStudy (anonymous):

the form of a quadratic is \[at ^{2}+bt+c\]

OpenStudy (anonymous):

quadratic equation \[-b \pm \frac{ \sqrt{b ^{2}-4ac} }{ 2a }\] the solution restriction is \[b ^{2}-4ac \ge 0\] since we only want one we can say \[b ^{2}-a(c-d) = 0\] so we can say that our solution is \[t = \frac{ -b }{ 2a }\] where \[t \ge 0\]

OpenStudy (anonymous):

thank you